Given open and closed pipes of the same length, which would have the lowest natural frequency:

A). The closed pipe

B). The open pipe

C). They both would have the same low frequency
Answer #1

Expressions for natural frequencies of open pipe and closed pipes are

fn = nv/2L for an open pipe,

and

fn = (2n - 1)v/4L for a closed pipe

For lowest frequency n =1,

open pipe frequency, f1 = v/2L and closed pipe frequency, f1 = v/4L .

As denominator in closed pipe equation is greater, it has lowest natural frequency.

Thus, the correct option is ( A )